### v1

# Embodied Intelligence for 3D Understanding: A Survey on 3D Scene Question Answering

Abstract
3D Scene Question Answering (3D SQA) represents an interdisciplinary task that integrates 3D visual perception and natural language processing, empowering intelligent agents to comprehend and interact with complex 3D environments. Recent advances in large multimodal modelling have driven the creation of diverse datasets and spurred the development of instruction-tuning and zero-shot methods for 3D SQA. However, this rapid progress introduces challenges, particularly in achieving unified analysis and comparison across datasets and baselines. This paper presents the first comprehensive survey of 3D SQA, systematically reviewing datasets, methodologies, and evaluation metrics while highlighting critical challenges and future opportunities in dataset standardization, multimodal fusion, and task design.
1 Introduction
Visual Question Answering (VQA) expands the scope of traditional text-based question answering Rajpurkar et al. (2016) by incorporating visual content, enabling the interpretation of images Antol et al. (2015), charts Masry et al. (2022), and documents Ding et al. (2024) to deliver context-aware responses. This capability facilitates a broader range of applications, including medical diagnostics Wu et al. (2022), financial analysis Xue et al. (2024), and assistance in academic research. Nevertheless, the growing demand of immersive 3D environments calls for even more natural and interactive question-answering systems. 3D Scene Question Answering (3D SQA) Azuma et al. (2022); Ye et al. (2021) addresses this by bridging visual perception He et al. (2016, 2017), spatial reasoning Guo et al. (2020), and language understanding in 3D environments Linghu et al. (2024), see Figure 1. Unlike traditional 3D tasks focused on object detection Qi et al. (2019); Li et al. (2023c) or segmentation Qi et al. (2017); Zou et al. (2024); He et al. (2025), 3D SQA integrates multimodal data, e.g., visual inputs and textual queries, to enable embodied systems capable of complex reasoning Szymanska et al. (2024). By leveraging spatial relationships, object interactions, and hierarchical scene structures within dynamic 3D environments, 3D SQA advances robotics, augmented reality, and autonomous navigation Huang et al. (2023), pushing the boundaries of multimodal AI and its potential in complex, real-world scenarios.
Figure 1: 2D Scene VQA and 3D SQA tasks. 3D SQA handles non-embodied as well as embodied tasks involving agent interactions within 3D scenes.
Early developments in 3D SQA were driven by manually annotated datasets like ScanQA Azuma et al. (2022) and SQA Ye et al. (2021), which aligned 3D point clouds with textual queries. Recently, programmatic generation methods, such as those used in 3DVQA Etesam et al. (2022) and MSQA Linghu et al. (2024), have enabled the creation of larger datasets with richer question types. The integration of Large Vision-Language Models (LVLMs) has further automated data annotation, leading to the development of more comprehensive datasets like LEO Huang et al. (2023) and Spartun3D Zhang et al. (2024b).
Methodologies have evolved alongside datasets, transitioning from closed-set approaches to LVLM-enabled techniques. Early methods Azuma et al. (2022); Ye et al. (2021) employed custom architectures combining point cloud encoders, e.g., PointNet++ Qi et al. (2017), and text encoders, e.g., BERT Kenton and Toutanova (2019), with attention-based fusion modules. However, they were constrained by predefined answer sets. The recent LVLM-based methods employ instruction-tuning Hong et al. (2023b); Huang et al. (2023) or zero-shot technique Yin et al. (2024); Linghu et al. (2024) while adapting models like GPT-4 Achiam et al. (2023), which reduces dependence on task-specific annotations. However, these methods also face challenges in ensuring dataset quality and addressing evaluation inconsistencies.
To analyse the emerging challenges in 3D SQA and facilitate their systematic handling, this paper provides the first comprehensive survey of this research direction. We focus on three fundamental aspects of this area, namely; (i) the objectives of 3D SQA, (ii) datasets needed to support these objectives, and (iii) models being developed to achieve these objectives. We review the evolution of datasets and methodologies, highlighting trends in the literature, such as the shift from manual annotation to LVLM-assisted generation, and the progression from closed-set to zero-shot methods. Additionally, we discuss challenges in multimodal alignment and evaluation standardization, offering insights into the future direction of the field. The paper outline that follows an organized structure of the existing 3D SQA literature, is provided in Figure 3 in the appendix.
Notation Definition $S$ A 3D scene representation. $S^{(p)}$ Point cloud representation of the scene: $S^{(p)}=\{(x_{i},y_{i},z_{i})\mid i=1,\dots,N\}$, where $x_{i},y_{i},z_{i}\in\mathbb{R}$ are coordinates. $S^{(m)}$ A set of multi-view images: $S^{(m)}=\{I_{1},I_{2},\dots,I_{K}\}$. $Q$ Multimodal query. $Q^{(t)}$ Textual query: $Q^{(t)}=(w_{1},w_{2},\dots,w_{L})$, where $w_{i}$ is a textual token. $Q^{(e)}$ Egocentric images in the query. $Q^{(o)}$ Object-level point clouds: $Q^{(o)}=\{(x_{j},y_{j},z_{j})\mid j=1,\dots,M\}$, $x_{j},y_{j},z_{j}\in\mathbb{R}$. $T$ A textual answer: $T=(t_{1},t_{2},\dots,t_{R})$. $B$ A set of 3D bounding boxes for objects referenced in the answer: $B=\{b_{1},b_{2},\dots,b_{M}\}$. $\mathcal{F}$ The task function mapping $(S,Q)\mapsto(T,B)$.
Table 1: 3D SQA task notations.
2 Preliminaries
The 3D SQA task involves comprehending a 3D scene $S$ and a query $Q$ to produce a textual answer $T$ and, optionally, spatial information $B$, such as bounding boxes for relevant objects. The 3D scene can be represented using modalities like point clouds $S^{(p)}$, multi-view images $S^{(m)}$, or their combinations, while the query may include textual input $Q^{(t)}$, egocentric images $Q^{(e)}$, or object-level point clouds $Q^{(o)}$. The task is formally defined as $\mathcal{F}:(S,Q)\mapsto(T,B)$, bridging multimodal reasoning and spatial understanding for comprehensive 3D scene analysis. For more details on this formulation, we refer to Table 1.
3 Datasets
Dataset Source Scene Q&A Collection Modality Suited Grounding
ScanQA Azuma et al. (2022) SCN Dai et al. (2017) 800 41K Template $S^{(p)}$$\boldsymbol{\times}$
SQA Ye et al. (2021) SCN Dai et al. (2017) 800 6K Human $S^{(p)}$$\boldsymbol{\times}$$\boldsymbol{\times}$
FE-3DGQA Zhao et al. (2022) SCN Dai et al. (2017) 703 20K Human $S^{(p)}$$\boldsymbol{\times}$
CLEVR3D Yan et al. (2023) 3RS Wald et al. (2019) 8,771 60K Template $S^{(p)}$$\boldsymbol{\times}$$\boldsymbol{\times}$
3DVQA Etesam et al. (2022) SCN Dai et al. (2017) 707 500K Template $S^{(p)}$$\boldsymbol{\times}$$\boldsymbol{\times}$
SQA3D Ma et al. (2022) SCN Dai et al. (2017) 650 33.4K Human $S^{(p)}$$\boldsymbol{\times}$
ScanScribe Zhu et al. (2023) SR Chen et al. (2020)+R3D Achlioptas et al. (2020) 2,995 56K LLM-assisted $S^{(p)}$$\boldsymbol{\times}$
3DMV-VQA Hong et al. (2023a) HM3d Ramakrishnan et al. (2021) 5K 50K Template $S^{(m)}$$\boldsymbol{\times}$$\boldsymbol{\times}$
OpenEQA Majumdar et al. (2024) SCN, HM3d Ramakrishnan et al. (2021); Yadav et al. (2023) 180 1.6K Human $S^{(m)}$$\boldsymbol{\times}$$\boldsymbol{\times}$
Spartun3D Zhang et al. (2024b) 3RS Wald et al. (2019) - 123K LLM-assisted $\{S^{(m)},S^{(p)}\}$$\boldsymbol{\times}$
MSQA Linghu et al. (2024) SCN, 3RS, ARK Baruch et al. (2021) - 254K LLM-assisted $\{S^{(m)},S^{(p)}\}$$\boldsymbol{\times}$
LEO Huang et al. (2023) SCN+3RS Wald et al. (2019) 3K 83K LLM-assisted $\{S^{(m)},S^{(p)}\}$
M3DBench Li et al. (2023b) ScanQA Azuma et al. (2022) - 320K LLM-assisted $\{S^{(m)},S^{(p)}\}$
3D-LLM Hong et al. (2023b) Objaverse Deitke et al. (2023) - 300K LLM-assisted $\{S^{(m)},S^{(p)}\}$
LAMM Yin et al. (2024) - - 186K LLM-assisted $\{S^{(m)},S^{(p)}\}$
Table 2: Comparison of 3D SQA datasets. Source abbreviations: SCN = ScanNet, 3RS = 3RScan, HME = HM3D, ARK = ARKitScenes, SR+R3D = ScanRefer + ReferIt3D. Modality abbreviations: $S^{(p)}$ = Point Cloud, $S^{(v)}$ = Video, $S^{(m)}$ = Multi-view Images. Suited: Indicates if the dataset is an Embodied 3D SQA dataset, requiring the agent to consider its state when answering.
Importance of datasets for contemporary 3D SQA can not be overemphasized. Existing datasets vary widely in scene representation, scale, and query complexity. To provide a systematic overview of the existing datasets, this section is organized as two main parts: Dataset Structure, which explores the representation and scale of scenes and queries. and QA Pair Creation, which examines methodologies for generating question-answer pairs.
3.1 Dataset Structure
In the data-driven domain of 3D SQA, structure of datasets significantly influences the scope of the tasks they support. Current datasets differ widely in their representations of 3D scenes, encompassing point clouds, multi-view images, and egocentric perspectives, as well as in the formats of their queries, which range from basic textual inputs to complex multimodal, embodied descriptions. Key dataset attributes such as scale, diversity of modalities, and query complexity significantly influence the design requirements and performance capabilities of 3D SQA models. Table 2 summarizes the key features of existing real-world 3D SQA datasets, providing an overview of their scene representations, query modalities, and scales. In Figure 2, we illustrate the typical dataset generation workflow at a higher level of abstraction.
3.1.1 Scene Modalities and Scale
Broadly, the development of 3D SQA datasets has progressed along a timeline evolving from synthetic environments to realistic 3D representations.
Synthetic 3D Datasets: The development of 3D SQA began with pseudo-3D datasets that utilized synthetic environments to simulate scene-level QA tasks. For example, EmbodiedQA Das et al. (2018) generated the dataset by selecting real scenes from the SUNCG Song et al. (2017) subset within the House3D Wu et al. (2018) simulator. These datasets were validated by human annotators to ensure quality. IQA Gordon et al. (2018) expanded this effort by introducing the IQUAD V1 dataset with 75,000 questions paired with unique scene configurations, leveraging the AI2-THOR Kolve et al. (2017) environment. MP3D-EQA Wijmans et al. (2019) and MT-EQA Yu et al. (2019) further incorporated depth maps and multi-target QA tasks, respectively, while remaining confined to synthetic SUNCG Song et al. (2017) scenes.
Point Cloud Datasets: The transition to real-world 3D SQA tasks was marked by the introduction of datasets based on 3D point clouds Rusu and Cousins (2011). ScanQA Azuma et al. (2022) and SQA Ye et al. (2021) established foundational benchmarks for this direction. Both datasets were constructed using ScanNet Dai et al. (2017), with ScanQA generating 41K QA pairs across 800 scenes, and SQA providing 6K manually curated QA pairs with higher linguistic accuracy. Building on these efforts, FE-3DGQA Zhao et al. (2022) selected 703 specific scenes from ScanNet and annotated 20K QA pairs, emphasizing foundational QA tasks with dense bounding box annotations to enable spatial grounding. CLEVR3D Johnson et al. (2017) utilized functional programs and text templates to generate four times the number of questions in ScanQA, introducing a broader range of attributes and question types. Subsequently, 3DVQA Etesam et al. (2022) expanded on CLEVR3D’s framework, leveraging 3D semantic scene graphs and template-based pipelines to generate questions and answers. By selecting 707 scenes, 3DVQA produced 500K QA pairs, significantly enriching task diversity and complexity. Similarly, SQA3D Ma et al. (2022) curated 33.4K manually annotated QA pairs across 650 scenes, focusing on linking queries to agent position and orientation.
Multi-View Datasets: To better align with human perception, multi-view datasets have been introduced, focusing on reasoning across different perspectives rather than relying solely on single point cloud representations. In this direction, 3DMV-VQA Hong et al. (2023a) includes 5K scenes from the HM3D dataset Ramakrishnan et al. (2021), generating 50K QA pairs. The images are rendered using the Habitat framework Ramakrishnan et al. (2021); Savva et al. (2019); Szot et al. (2021), emphasizing multi-view reasoning. On the other hand, OpenEQA Majumdar et al. (2024) not only selects scenes from HM3D but also incorporates Gibson Xia et al. (2018) and ScanNet Dai et al. (2017), ultimately choosing 180 high-quality scenes with 1.6K QA pairs. Unlike other datasets, it prioritizes quality over scale, making it a significant contribution to high-quality 3D QA benchmarks.
Multimodal Datasets: Recent advances in 3D SQA datasets emphasize integrating point clouds, images, and textual data to form rich multimodal representations. These approaches aim to capture spatial, semantic, and contextual cues for more comprehensive scene understanding. A notable example is Spartun3D Zhang et al. (2024b), which selects scenes from 3RScan Wald et al. (2019) and generates 123K QA pairs focused on situational tasks. Similarly, MSQA Linghu et al. (2024) builds 254K QA pairs from multimodal datasets Dai et al. (2017); Wald et al. (2019); Baruch et al. (2021), using point clouds and object images as inputs to better align with real-world embodied intelligence scenarios.
With the popularity of LLMs, instruction tuning datasets have also emerged as an important extension of multimodal datasets, enhancing the generalization capabilities of 3D SQA models by aligning 3D data with textual descriptions. For instance, ScanScribe Zhu et al. (2023) collects RGB-D scans of indoor scenes from ScanNet and 3R-Scan, incorporating diverse object instances from Objaverse Deitke et al. (2023). It uses QA pairs from ScanQA and referential expressions from ScanRefer Chen et al. (2020) and ReferIt3D Achlioptas et al. (2020), generating 56.1K object instances from 2,995 scenes through templates and GPT-3 Brown (2020). Similarly, LEO Huang et al. (2023) constructs 83K 3D-text pairs by collecting captions at object, object-in-scene, and scene levels Luo et al. (2024); Achlioptas et al. (2020); Zhu et al. (2023); Chen et al. (2021).
Along similar lines, M3DBench Li et al. (2023b) leverages multiple existing and LLMs to generate 320K instruction-response pairs, enriching multimodal 3D data for a wide range of 3D-language tasks. 3D-LLM Hong et al. (2023b) creates over 300K 3D-text pairs using assets like Objaverse, ScanNet, and HM3D, while LAMM Yin et al. (2024) employs GPT-API and self-instruction methods Wang et al. (2022) to produce 186K language-image pairs and 10K language-3D pairs.
3.1.2 Query Modalities and Complexity
Figure 2: Dataset generation workflow.
In 3D SQA, a query represents the input question or prompt that, when paired with a 3D scene, guides the task of providing an answer. Over time, query modalities in 3D SQA have evolved from simple text-based inputs to more complex, multimodal, and agent-centric formats. Here, we summarise the datasets from the query modality perspective, which is a critical consideration for dataset selection in performance evaluation.
Basic Text Queries: Early 3D SQA datasets primarily employed straightforward text-based queries that focused on scene-level attributes, such as object counting or identification. These datasets aimed to evaluate foundational 3D scene understanding, often without considering the agent’s position, interaction, or perspective within the environment. For example, datasets like ScanQA Azuma et al. (2022) and SQA Ye et al. (2021) feature questions such as "How many chairs are in the room?". Such purely textual questions fail to capture complex embodied scenarios as they lack description of an agent’s spatial or contextual relationship with the scene. Consequently, these datasets are limited in scope, as reflected in Table 2, where the lack of Suited queries indicates their omission of agent-centric contexts. This limitation underscores the evolution toward richer, more contextualized datasets in the later 3D SQA research.
Agent-Centric Text Queries: The introduction of agent-centric descriptions marked a significant shift in query complexity. SQA3D Ma et al. (2022) was one of the first datasets to incorporate contextualized questions, where textual queries were enhanced with references to the agent’s position or orientation. In this case, a typical query might describe the agent’s location, such as "Sitting at the edge of the bed and facing the couch.". We mark datasets enabling such queries as Suited in Table 2.
Multimodal Agent-Centric Queries: Recently, SPARTUN3D Zhang et al. (2024b) and MSQA Linghu et al. (2024) introduced richer spatial descriptions and multimodal query inputs. The former provided detailed spatial information, enabling queries such as "You are standing beside a trash bin while there is a toilet in front of you.". Similarly, MSQA integrated textual descriptions, explicit spatial coordinates, and agent orientation in the queries. Additionally, first-person view images were included. These multimodal approaches enable more realistic scenarios by combining spatial, visual, and linguistic contexts.
Instruction-Tuned Queries: Recent datasets, such as ScanScribe Zhu et al. (2023), LEO Huang et al. (2023), and M3DBench Li et al. (2023b), have also expanded query modalities further to support instruction tuning tasks. They leverage agent-centric queries enriched with multimodal inputs, such as spatially grounded textual descriptions and multimodal instructions. For example, LEO incorporates multimodal instructions to fine-tune models for agent tasks like real-time navigation or object interaction. M3DBench focuses on generalization across diverse real-world tasks by utilizing rich multimodal data. These instruction-tuning datasets ensure models are well-equipped to address practical, real-world tasks by aligning textual instructions with spatial and visual contexts.
3.2 QA Pair Creation
The creation of question-answer (QA) pairs defines the scope and complexity of 3D SQA tasks. Early datasets relied on manual annotation, while recent efforts have adopted templates and LVLMs to improve scalability and diversity. These advances have enabled datasets to include a wider range of question types, from object identification to spatial relationships and task-specific queries.
3.2.1 Methods for QA Pair Generation
QA pair generation in 3D SQA datasets balances between manual annotation, template-based pipelines, and LLM-assisted methods. Manual annotation ensures high-quality and contextual accuracy, while template-based approaches enable scalable generation with logical consistency. Recently, LLMs have further automated the process, enabling diverse multimodal QA pairs at scale. This progression, also apparent in Figure 2, reflects the evolution of dataset creation techniques.
Template-Based Generation: Template-based generation was introduced as an early solution for scalable QA pair creation. ScanQA Azuma et al. (2022) exemplified this approach by utilizing a T5-based QA generation model Raffel et al. (2020) to generate seed questions from ScanRefer Chen et al. (2020). Similarly, datasets like CLEVR3D Yan et al. (2023), 3DVQA Etesam et al. (2022), and 3DMV-VQA Hong et al. (2023a) leveraged 3D Semantic Scene Graphs to programmatically generate diverse and logically consistent QA pairs, improving scalability and task diversity. While the template-based approach enables large-scale datasets, the generated questions often lack contextual specificity and may sometimes result in overly generic queries.
Manual Annotation: Researchers have also pursued manual annotation to address the limitations of template-based methods. Manual approaches prioritize linguistic precision and contextual relevance, creating datasets that are smaller in scale but of higher quality. For instance, SQA Ye et al. (2021) curated 6K QA pairs with an emphasis on linguistic accuracy, while FE-3DGQA Zhao et al. (2022) selected 703 scenes from ScanNet Dai et al. (2017) and annotated 20K QA pairs, grounding answers with bounding box annotations. Similarly, OpenEQA Majumdar et al. (2024) curated 1.6K QA pairs from 180 high-quality scenes. SQA3D Ma et al. (2022) contributed 33.4K QA pairs across 650 scenes, tailored specifically for agent-centric tasks. Despite their time-intensive nature, fully curated datasets play a critical role in ensuring accuracy and contextual alignment, complementing the template-based methods.
LLM-Assisted Generation: Recent methods have increasingly leveraged LLMs to automate the generation of QA pairs, enhancing both scalability and diversity. Notable examples include Spartun3D Zhang et al. (2024b) and MSQA Linghu et al. (2024), both of which utilize scene graphs to structure spatial and semantic relationships. Spartun3D employs GPT-3.5 to generate agent-centric questions, emphasizing situated reasoning and exploration, resulting in 123K QA pairs. MSQA takes a similar approach with GPT-4V, focusing on situated QA generation guided by semantic scene graphs, producing 254K QA pairs.
Additionally, LLMs have been instrumental in constructing instruction tuning datasets to improve model generalization across diverse multimodal tasks. ScanScribe Zhu et al. (2023) utilizes GPT-3 to transform ScanRefer annotations into scene descriptions using template-based refinement. LEO Huang et al. (2023) adopts GPT-4 with Object-centric Chain-of-Thought (O-CoT) prompting to ensure logical consistency. M3DBench Li et al. (2023b) and 3D-LLM Hong et al. (2023b) use GPT-4 to create multimodal prompts based on object attributes and scene-level inputs. Together, these datasets demonstrate the growing role of LLMs in automating the generation of high-quality, multimodal data for 3D SQA.
3.2.2 Question Design in 3D SQA
With advancements in language and vision modelling, 3D SQA questions have evolved along several dimensions: from simple to complex tasks, non-situated to situated contexts, and static to dynamic scenarios. To exemplify the nature of these questions, we enlist the common 3D SQA tasks and representative question in Table 6 in the appendix.
Task Complexity - From Basic to Advanced: 3D SQA covers a diverse spectrum of question tasks designed to assess models’ understanding of 3D environments and their reasoning abilities. Basic tasks, such as object identification, spatial reasoning, attribute querying, object counting, and attribute comparison, are featured in datasets like SQA Ye et al. (2021), ScanQA Azuma et al. (2022), FE-3DGQA Zhao et al. (2022), 3DVQA Etesam et al. (2022) and CLEVR3D Yan et al. (2023). Among these, FE-3DGQA introduced more complex, free-form questions that require models not only to ground answer-relevant objects but also to identify contextual relationships between them. Similarly, CLEVR3D emphasized relational reasoning by incorporating questions that integrate objects, attributes, and their interrelationships, pushing models further to handle intricate contextual dependencies.
As 3D SQA evolves, tasks demanding a deeper understanding of spatial and visual context have emerged, challenging models to engage with dynamic and context-aware reasoning. These tasks include multi-hop reasoning (SQA3D Ma et al. (2022)), navigation (SQA3D Ma et al. (2022), LEO Huang et al. (2023), 3D-LLM Hong et al. (2023b), M3DBench Li et al. (2023b), MSQA Linghu et al. (2024)), robotic manipulation (LEO), object affordance (Spartun3D Zhang et al. (2024b)), functional reasoning (OpenEQA Majumdar et al. (2024)), multi-round dialogue (LEO, M3DBench, 3D-LLM), planning (LEO, M3DBench, Spartun3D), and task decomposition (3D-LLM). These advanced tasks challenge models to dynamically reason and navigate complex 3D environments while capturing intricate spatial and relational details. Notably, OpenEQA Majumdar et al. (2024) stands out as the first open-vocabulary dataset for embodied question answering.
Situated vs. Non-Situated Questions: Based on the required level of interaction and contextual understanding, 3D VQA questions can be categorized into situated and non-situation types. The latter focus on static reasoning, testing a model’s ability to interpret spatial relationships, attributes, and object properties within fixed 3D scenes. Datasets like SQA Ye et al. (2021), ScanQA Azuma et al. (2022), FE-3DGQA Zhao et al. (2022), 3DVQA Etesam et al. (2022), CLEVR3D Yan et al. (2023), and LAMM Yin et al. (2024) primarily include non-situated questions that evaluate understanding within static spatial contexts.
Conversely, situated questions involve dynamic reasoning, requiring interaction with the 3D environment and comprehension of contextual or sequential information. These questions test models’ ability to navigate, plan, and adapt to dynamic scenarios and often include temporal or embodied elements. Situated questions appear in datasets like SQA3D Ma et al. (2022), LEO Huang et al. (2023), 3D-LLM Hong et al. (2023b), M3DBench Li et al. (2023b), MSQA Linghu et al. (2024), Spartun3D Zhang et al. (2024b), 3DMV-VQA Hong et al. (2023a), and OpenEQA Majumdar et al. (2024). This categorization enables a comprehensive evaluation of 3D VQA systems.
Temporal Aspect in 3D SQA: Most 3D SQA datasets limit questions to a single time slot, reflecting the static nature of the environments they evaluate. This restriction simplifies reasoning by focusing on a specific moment within the 3D scene. However, datasets like OpenEQA Majumdar et al. (2024) now introduce dynamic scenarios that allow for multiple time slots, enabling tasks that require episodic memory and active exploration. This temporal dimension challenges models to integrate sequential information and represents a significant step forward for advancing 3D SQA.
3.3 Evaluating LLM-Generated 3D Datasets
While LLM adoption has significantly advanced 3D SQA datasets, ensuring their quality, reliability, and practical utility remains an open challenge. Current evaluation methods primarily rely on manual assessments. For example, LEO Huang et al. (2023) evaluates QA pairs through expert review, reporting metrics like overall accuracy and contextual relevance. MSQA Linghu et al. (2024) adopts a comparative approach, sampling QA pairs from its dataset and comparing them against a benchmark dataset such as SQA3D Ma et al. (2022), with scores based on contextual accuracy, factual correctness, and overall quality. Similarly, Spartun3D Zhang et al. (2024b) employs expert validation by randomly sampling instances to ensure that the generated data meets expected quality standards. These manual evaluations provide valuable insights into dataset quality but face limitations in scalability, labour intensity, and subjectivity.
To address these limitations, automated evaluation frameworks are currently needed. Potential solutions include embedding-based metrics for semantic alignment, logical consistency checks for QA coherence, and task-specific metrics for spatial accuracy and multimodal integration.
4 Evaluation Metrics
Standardized evaluation metrics are crucial to gauge advances in 3D SQA and ensure dataset suitability for downstream tasks. Contemporary 3D SQA literature either uses traditional or LLM-based metrics for the evaluation purpose.
Traditional metrics: 3D SQA methods often employ quantitative measures of linguistic relevance and correctness for evaluation. Commonly used metrics include Exact Match (EM@1, EM@10), which assesses whether the generated answers match ground truth exactly, and language generation metrics such as BLEU Papineni et al. (2002), ROUGE-L Lin (2004), METEOR Banerjee and Lavie (2005), CIDEr Vedantam et al. (2015), and SPICE Anderson et al. (2016). These metrics were initially employed by ScanQA Azuma et al. (2022) and have since been used for datasets like CLEVR3D Yan et al. (2023), 3DGQA Zhao et al. (2022), and ScanScribe Zhu et al. (2023). While effective for evaluating linguistic accuracy and diversity, traditional metrics are generally limited in capturing the nuanced reasoning and contextual understanding required for 3D SQA tasks.
LLM-based metrics: The emerging evaluation paradigm in 3D SQA employs LLM-based metrics, leveraging the reasoning capabilities of models like GPT. For instance, OpenEQA Majumdar et al. (2024) employs GPT to evaluate the contextual relevance and correctness of generated answers, introducing a metric that eventually computes the Mean Relevance score. Similarly, MSQA Linghu et al. (2024) uses GPT to assess the quality of answers based on nuanced reasoning, aligning them with contextual expectations. Compared to traditional metrics, LLM-based methods currently excel at simulating real-world reasoning and capturing semantic subtleties, making them particularly valuable for evaluating complex multimodal tasks.
In summary, traditional metrics provide a strong foundation for evaluating linguistic and structural quality, while LLM-based metrics offer deeper insights into contextual alignment and reasoning. Combining the complementary properties of these metrics can offer a comprehensive framework for assessing 3D SQA performance.
Method Type Scene Modality Scene Encoder Text Encoder Answer Module
ScanQA Azuma et al. (2022) T-S $S^{(p)}$ VoteNet Qi et al. (2019) BiLSTM Graves and Graves (2012) MLP
3DQA-TR Ye et al. (2021) T-S $S^{(p)}$ Group-Free Liu et al. (2021b) BERT Kenton and Toutanova (2019) MLP
TransVQA3D Yan et al. (2023) T-S $S^{(p)}$ PointNet++ Qi et al. (2017) BERT Kenton and Toutanova (2019) MLP
FE-3DGQA Zhao et al. (2022) T-S $S^{(p)}$ PointNet++ Qi et al. (2017) T5 Raffel et al. (2020) Linear Layer
SIG3D Man et al. (2024) T-S $S^{(p)}$ OpenScene Peng et al. (2023) BiLSTM Graves and Graves (2012) MLP
3D-CLR Hong et al. (2023a) T-S $S^{(m)}$ CLIP-LSeg Li et al. (2022) CLIP Radford et al. (2021) 3D CNN
BridgeQA Mo and Liu (2024) T-S $\{S^{(m)},S^{(p)}\}$ VoteNet&BLIP Li et al. (2023a) BLIP Li et al. (2023a) Transformer
3DVLP Zhang et al. (2024a) P-B $S^{(p)}$ PointNet++ Qi et al. (2017) CLIP Radford et al. (2021) MLP
CLIP-Guided Parelli et al. (2023) P-B $S^{(p)}$ VoteNet&Transformer CLIP Radford et al. (2021) MLp
Multi-CLIP Delitzas et al. (2023) P-B $S^{(p)}$ VoteNet&Transformer CLIP Radford et al. (2021) MLP
3D-VisTA Zhu et al. (2023) P-B $S^{(p)}$ PointNet++ Qi et al. (2017) BERT MLP
GPS Jia et al. (2025) P-B $S^{(p)}$ PointNet++ Qi et al. (2017) Transformer Vaswani (2017) Transformer
LM4Vision Pang et al. (2023) P-B(w I-T) $S^{(p)}$ VoteNet Qi et al. (2019) LSTM Hochreiter (1997) LLaMA Touvron et al. (2023)
3D-LLM Hong et al. (2023b) P-B(w I-T) $S^{(m)}$ BLIP2 Li et al. (2023a) BLIP2 Li et al. (2023a) BLIP2 Li et al. (2023a)
LEO Huang et al. (2023) P-B(w I-T) $S^{(p)}$ PointNet++& ST Chen et al. (2022) ConvNext Liu et al. (2022) Vicuna Chiang et al. (2023)
LAMM Yin et al. (2024) P-B(w I-T) $S^{(p)}$ PointNet++ Qi et al. (2017) SentencePiece Kudo (2018) Vicuna Chiang et al. (2023)
M3DBench Li et al. (2023b) P-B(w I-T) $S^{(p)}$ PointNet++& Transformer Opt Zhang et al. (2022) Opt Zhang et al. (2022)
SQA3D Ma et al. (2022) Z-S $S^{(p)}$ Scan2Cap Chen et al. (2021) GPT-3 Brown (2020) GPT-3
LAMM Yin et al. (2024) Z-S $S^{(p)}$ PointNet++ Qi et al. (2017) SentencePiece Kudo (2018) Vicuna
EZSG Singh et al. (2024) Z-S $S^{(m)}$ GPT-4V Yang et al. (2023) GPT-4V Yang et al. (2023) GPT-4V
OpenEQA Majumdar et al. (2024) Z-S $S^{(m)}$ GPT-4V Yang et al. (2023) GPT-4V Yang et al. (2023) GPT-4V
MSQA Linghu et al. (2024) Z-S $S^{(m)}$ GPT-4o Achiam et al. (2023) GPT-4o Achiam et al. (2023) GPT-4o
LEO Huang et al. (2023) Z-S $\{S^{(m)},S^{(p)}\}$ PointNet++& ST Chen et al. (2022) ConvNext Liu et al. (2022) Vicuna Chiang et al. (2023)
Spartun3D-LLM Zhang et al. (2024b) Z-S $\{S^{(m)},S^{(p)}\}$ PointNet++ Qi et al. (2017) CLIP Radford et al. (2021) Vicuna
Table 3: Overview of techniques for 3D SQA. Methods are categorized as Task-Specific (T-S), Pretraining-Based (P-B) and Zero-Shot (Z-S). P-B (w I-T) denotes Pretraining-Based methods further enhanced with Instruction Tuning to better adapt to task-specific instructions. Scene modalities are represented as $S^{(p)}$ for Point Cloud, $S^{(m)}$ for Image, and $\{S^{(m)},S^{(p)}\}$ for Multimodal.
5 Taxonomy of 3D SQA Methods
3D SQA methods can be categorized into three primary types, as shown in Table 3. i) Task-Specific Methods rely on predefined answers and specialized architectures to address specific tasks. ii) Pretraining-Based Methods leverage large-scale datasets to align multimodal representations and fine-tune for task-specific objectives. iii) Zero-Shot Learning Methods also utilize pretrained LLMs and VLMs to generalize to new tasks, albeit without additional fine-tuning. These categories underpin the field’s evolution from task-specific to scalable approaches that harness the capabilities of advanced multimodal models, reflecting the increasing focus on flexibility and adaptability in 3D SQA systems. Figure 4 in the appendix illustrates the common high-level pipeline of the methods.
5.1 Task-Specific Methods
These methods are designed for specific tasks using closed-set classification approach.
Point Cloud Methods: 3D SQA methods for point clouds follow a modular pipeline of scene and query encoding, feature fusion, and answer prediction. Early approaches like ScanQA Azuma et al. (2022) employed VoteNet Qi et al. (2019) and PointNet++Qi et al. (2017) to extract spatial features, while textual queries were encoded using GloVe Pennington et al. (2014) and BiLSTM Graves and Graves (2012). Fusion was achieved through transformer-based modules.
Building on this foundation, later methods introduced more sophisticated encoders and fusion strategies. For example, 3DQA-TR Ye et al. (2021) replaced VoteNet with Group-Free Liu et al. (2021b) for finer-grained scene encoding and adopted BERT Kenton and Toutanova (2019) for query encoding. Fusion was further streamlined by directly integrating features via a text-to-3D transformer Ye et al. (2021), enabling more direct question-to-answer mappings. Similarly, TransVQA3D Yan et al. (2023) enhanced feature interaction by introducing SGAA for fusion, focusing on global and local semantics in scenes.
For the datasets requiring spatial grounding, FE-3DGQA Zhao et al. (2022) advanced the pipeline by using PointNet++ Qi et al. (2017) for spatial feature extraction and T5 Raffel et al. (2020) for textual encoding, complemented by an attention mechanism Zhao et al. (2021); Liu et al. (2021a) to align text with dense spatial annotations. The recently proposed SIG3D Man et al. (2024) focuses on context-aware tasks in embodied intelligence. It encodes scenes using voxel-based tokenization and employs anchor-based contextual estimation to determine the agent’s position and orientation.
Multi-view and 2D-3D Methods: A few methods also use multi-view images to enhance 3D SQA performance. For example, 3D-CLR Hong et al. (2023a) constructs compact 3D scene representations by leveraging multi-view images and optimizing 3D voxel grids. On the other hand, 2D-3D methods like BridgeQA Mo and Liu (2024) combine 2D image features from pretrained VLMs Radford et al. (2021); Li et al. (2023a) with 3D object-level features obtained through VoteNet Qi et al. (2019). Both feature types are aligned with text features encoded by the VLM’s text encoder and fused using a vision-language transformer, enabling free-form answers.
Advances in Text Encoders: The evolution of text encoders in 3D SQA reflects the increasing demands for contextual and multimodal understanding by the models. Early methods employed BiLSTM Graves and Graves (2012) and BERT Kenton and Toutanova (2019) for basic semantic and syntactic feature extraction, as seen in ScanQA Azuma et al. (2022) and 3DQA-TR Ye et al. (2021). More recent approaches, such as FE-3DGQA Zhao et al. (2022), leverage transformer-based models like T5 Raffel et al. (2020) for richer linguistic embeddings. Meanwhile, multimodal models like CLIP Radford et al. (2021) in 3D-CLR Hong et al. (2023a) and BLIP Li et al. (2023a) in BridgeQA Mo and Liu (2024) have been instrumental in aligning textual and visual features. These advancements highlight a shift towards models that seamlessly integrate text with 3D spatial representations for improved performance. Task-specific methods are typically evaluated on the ScanQA and SQA3D datasets. Tables 4 and 5 in the appendix provide performance comparison summaries on these dataset for existing methods.
5.2 Pretraining-Based Methods
Pretraining-based approaches in 3D SQA have transitioned from traditional methods that emphasize explicit alignment of spatial and textual embeddings to instruction-tuning paradigms that harness large pretrained models. These methods strike a balance between task-specific adaptation and generalization to address challenges of scalability.
Traditional Pretraining Methods: These methods focus on aligning 3D spatial features with rich 2D visual and linguistic representations. Parelli et al. (2023) utilized a trainable 3D scene encoder based on VoteNet Qi et al. (2019) to extract object-level features, which are further refined using a Transformer layer to model inter-object relationships. Multi-CLIP Delitzas et al. (2023) introduces multi-view rendering and robust contrastive learning to enhance the integration of 3D spatial features with 2D representations. Zhang et al. (2024a) introduced object-level cross-contrastive and self-contrastive learning tasks during pretraining to improve cross-modal alignment. Jia et al. (2025) adopted a hierarchical contrastive alignment strategy, combining object-level, scene-level, and referential embeddings to enhance cross-modal and intra-modal feature integration.
Diverging from these contrastive learning approaches, 3D-VisTA Zhu et al. (2023) employs a unified Transformer-based framework Vaswani (2017) to align 3D scene features with textual representations. Instead of relying on extensive annotations, it leverages self-supervised objectives to optimize multimodal alignment He et al. (2021); Radford et al. (2019). This shift from task-specific pretraining to self-supervised learning is a noteworthy development for efficient and robust 3D SQA.
Instruction-Tuning Methods: Pretrained foundation models learn general geometric and semantic representations from large-scale unsupervised data at high computational cost. Instruction-tuning methods exploit the generalization abilities of these models by leveraging pretrained LLMs or VLMs as frozen encoders. These methods retain the parameters of the encoders, making minimal modifications, typically through lightweight task-specific layers, to adapt to downstream tasks. Recent approaches, such as LM4Vision Pang et al. (2023), 3D-LLM Hong et al. (2023b), LEO Huang et al. (2023), M3DBench Li et al. (2023b), and LAMM Yin et al. (2024), exemplify this shift.
LM4Vision Pang et al. (2023) employs a frozen LLaMA Touvron et al. (2023) encoder and trains lightweight task-specific layers for alignment with the 3D QA tasks. Similarly, 3D-LLM builds upon the BLIP2 Li et al. (2023a), adding a task-specific head while keeping the base model frozen. In contrast, LEO, M3DBench, and LAMM utilize Vicuna Chiang et al. (2023), a derivative of LLaMA, to integrate textual and multimodal inputs. LEO incorporates object-centric and scene-level captions for enhanced multimodal reasoning. By leveraging the extensive knowledge encoded in LLMs or VLMs, these methods bypass the need for large task-specific pretraining datasets. Additionally, instruction-tuning methods are also effective in zero- and few-shot scenarios.
5.3 Zero-Shot Learning Methods
Zero-shot has emerged as a promising learning paradigm for 3D SQA, enabling models to infer answers to unseen tasks without task-specific fine-tuning. Current zero-shot 3D SQA methods can be broadly categorized into: text-driven, image-driven, and multimodal alignment approaches.
Text-Driven Approaches: These methods convert 3D scene information into textual descriptions, which are then used with a question in pretrained LLMs or VLMs for zero-shot inference. An example is SQA3D Ma et al. (2022), which uses Scan2Cap Chen et al. (2021) to generate scene descriptions and inputs them into GPT-3 Brown (2020) for answering questions. However, this approach overlooks the spatial structure of point clouds and images, limiting its ability to fully leverage 3D information. Similarly, LAMM Yin et al. (2024) extracts features from point clouds and text, but uses 3D data in a limited manner.
Image-Driven Approaches: These methods use VLMs to incorporate visual features like images or multi-view data along with text. For instance, MSQA Linghu et al. (2024) uses GPT-4o Achiam et al. (2023) with VLMs. Singh et al. (2024) tested unfinetuned GPT-4V Yang et al. (2023) on datasets like 3D-VQA and ScanQA Azuma et al. (2022), showing competitive performance in certain tasks. These methods are flexible and resource-efficient, but they still rely on text to represent spatial and object relationships, which is a potential limitation.
Multimodal Alignment Approaches: Techniques such as LEO Huang et al. (2023) and Spartun3D-LLM Zhang et al. (2024b), explicitly align visual and textual information during pretraining. LEO improves zero-shot performance by aligning object- and scene-level features, while Spartun3D-LLM employs an explicit module for aligning point clouds and text. These methods require relatively more training resources due to additional computations. Nevertheless, they offer an attractive trade-off between performance and efficiency.
Overall, in contemporary Zero-shot 3D SQA, Text-driven approaches are cost-effective and flexible but suffer from limited utilization of 3D data. Image-driven methods, which directly leverage VLMs for inference, also face limitations due to insufficient exploitation of 3D information. Multimodal alignment methods, while offering superior performance, have higher resource requirements.
6 Challenges and Future Directions
While 3D SQA has seen notable advancements, several critical challenges remain, limiting its potential for real-world applications. We outline key challenges and propose directions for future research.
Dataset Quality and Standardization. The rapid development of 3D SQA datasets in recent years has led to a fragmented landscape, with datasets varying widely in scope and modality. Integrating these datasets into unified benchmarks can offer the much needed standardised evaluation to catapult research in this direction. Additionally, while LLMs facilitate scalable dataset generation, they often introduce hallucinated information and contextual misalignments. Future research should focus on robust validation frameworks, leveraging human-in-the-loop systems or LLMs as validators.
Enhancing 3D Awareness in Zero-Shot. Current zero-shot models heavily rely on textual proxies, with limited utilization of 3D spatial and geometric features. Although multi-view approaches mitigate this issue to some extent, the lack of explicit 3D representation hampers their effectiveness for spatially complex tasks. Instruction-tuning methods face similar limitations. Future work needs to explore architectures that deeply integrate 3D features with linguistic and visual modalities to enhance generalization across diverse tasks. Additionally, an apparent direction for future research is to explore the balance between multimodal alignment and pretrained models in zero-shot 3D SQA to enhance both efficiency and performance.
Unified Evaluation. Absence of standardized and 3D SQA objective-specific evaluation metrics currently complicates meaningful evaluation and comparisons across datasets and models. Developing unified frameworks that incorporate multimodal metrics for spatial reasoning, contextual accuracy, and task-specific performance are currently required to enable accurate benchmarking and drive methodological innovation in 3D SQA.
Dynamic and Open-World Scenarios. Most existing methods and datasets focus on static, predefined environments, limiting applicability to real-world tasks. Future efforts need to emphasize more on dynamic, open-world settings, enabling models to handle real-time scene changes and novel queries. Incorporating embodied interactions, such as navigation and multi-step reasoning, will further align 3D SQA systems with real-world requirements.
Interpretable and Explainable 3D SQA Models. Current 3D SQA models often act as "black boxes", limiting their adoption in trust-critical domains like healthcare. Developing interpretable models that visualize 3D features, highlight relevant regions, or provide natural language explanations can enhance user trust and broaden their applicability.
Multimodal Interaction and Collaboration. 3D SQA systems are evolving toward more natural and interactive interfaces. Future research can explore integrating linguistic, gestural, and visual inputs to enable intuitive interaction with 3D scenes. Additionally, collaborative scenarios, such as architectural design or educational training, where multiple users interact with the system in real-time, offer a promising direction. Such systems could enhance communication and joint problem-solving, unlocking broader applications for 3D SQA.
Incorporating Temporal Dynamics. Most 3D SQA models currently ignore temporal dynamics of the scenes, whereas most of the real-world applications, such as traffic monitoring, robotic navigation, involve dynamic environments. Future research should aim to incorporate temporal dynamics into 3D SQA, allowing models to reason about scene changes over time. Leveraging temporal information, such as object movements, would enable these systems to better handle tasks requiring long-term temporal reasoning.
Model Efficiency and Deployment. Deploying 3D SQA systems on resource-constrained devices, such as mobile robots and edge AI agents, remains challenging due to high computational and memory demands. Future work should focus on lightweight architectures and optimization techniques, including pruning, quantization, and knowledge distillation, to enable efficient and real-time inference. Energy-efficient algorithms and scalable designs tailored for embedded systems will further enhance the practicality of 3D SQA in real-world applications.
By addressing these challenges, 3D SQA can advance toward robust, scalable, and versatile systems, driving significant progress in embodied intelligence and multimodal reasoning.
Appendix A Appendix
{forest}
for tree= forked edges, grow’=0, draw, rounded corners, node options=align=center, text width=2.2cm, s sep=5pt, calign=child edge, calign child=(n_children()+1)/2, [3D SQA Categories, text width=3cm, fill=gray!30, for tree=fill=brown!45 [Datasets, for tree=fill=blue!40 [Dataset Structure, for tree=fill=blue!20 [Scene Modalities and Scale, for tree=fill=blue!10 [Synthetic 3D Datasets] [Point Cloud Datasets] [Multi-View Datasets] [Multimodal Datasets] ] [Query Modalities and Complexity, for tree=fill=blue!10 [Basic Text Queries] [Agent-Centric Text Queries] [Multimodal Agent-Centric Queries] [Instruction-Tuned Queries] ] ] [QA Pair Creation, for tree=fill=blue!20 [Methods for QA Pair Generation [Template-Based Generation] [Manual Annotation] [LLM-Assisted Generation] ] [Question Design in 3D SQA [Task Complexity] [Situated vs. Non-Situated] [Temporal Aspec] ] ] [Evaluating LLM-Generated 3D Datasets, for tree=fill=blue!20] ] [Evaluation Metrics, for tree=fill=green!40 [Traditional metrics] [LLM-based metrics] ] [Methodological Taxonomy, for tree=fill=violet!40 [Task-Specific Methods, for tree=fill=violet!20 [Point Cloud Methods] [Multi-view and 2D-3D Methods] [Advances in Text Encoders] ] [Pretraining-Based Methods, for tree=fill=violet!20 [Traditional Pretraining Approaches] [Instruction-Tuning Methods] ] [Zero-Shot Learning Methods, for tree=fill=violet!20 [Text-Driven Approaches] [Image-Driven Approaches] [Multimodal Alignment Approaches] ] ] [Challenges and Future Directions, for tree=fill=red!40 [Dataset Quality ] [Zero-Shot] [Unified Evaluation] [Dynamic and Open-World, text width=4.2cm] [Interpretable and Explainable, text width=4.2cm] [Multimodal and Collaboration, text width=4.2cm] [Incorporating Temporal Dynamics, text width=4.2cm] ] ]
Figure 3: Graphical illustration of the hierarchical structure of 3D SQA literature adopted in this work. A systematic categorization is adopted for methodologies, datasets, and evaluation metrics.
Model EM@1 EM@10 B-1 B-2 B-3 B-4 R M C
ScanQA Azuma et al. (2022) 21.05 51.23 30.24 20.40 15.11 10.08 33.30 13.14 64.90
FE-3DGQA Zhao et al. (2022) 22.26 54.51 - - - - - - -
3DVLP Zhang et al. (2024a) 24.03 57.91 - - - - - - -
CLIP-Guided Parelli et al. (2023) 23.92 - 32.82 - - 14.64 35.15 13.94 69.53
Multi-CLIP Delitzas et al. (2023) 24.02 - 32.63 - - 12.65 35.46 13.97 68.70
LAMM Yin et al. (2024) - - - - - - - - -
3D-VisTA Zhu et al. (2023) 27.00 57.90 - - - 16.00 38.60 15.20 76.60
3D-LLM Hong et al. (2023b) 21.20 - 39.30 25.20 18.40 12.00 37.85 15.10 74.50
SceneVerse Jia et al. (2025) 22.70 - - - - - - - -
ESZG Singh et al. (2024) 18.01 18.01 30.24 20.40 15.11 10.08 33.33 13.14 64.86
SIG3D Man et al. (2024) - - 39.50 - - 12.40 35.90 13.40 68.80
Human 51.60 - - - - - - - -
Table 4: Performance comparison of existing models on ScanQA datasets. EM@1 and EM@10 refer to exact match accuracy for top-1 and top-10 answers, respectively. B-1 to B-4 represent BLEU-1 to BLEU-4 scores. R, M, and C stand for ROUGE, METEOR, and CIDEr metrics, respectively. Higher values are more desirable for all metrics.
Model What Is How Can Which Others Avg
ScanQA Azuma et al. (2022) 28.60 65.00 47.30 66.30 43.90 42.90 45.30
SQA3D Ma et al. (2022) 33.48 66.10 42.37 69.53 43.02 46.40 47.02
SQA3D (GPT-3) Ma et al. (2022) 39.67 45.99 40.47 45.56 36.08 38.42 41.00
Multi-CLIP Delitzas et al. (2023) - - - - - - 48.00
3D-VisTA Zhu et al. (2023) 34.80 63.30 45.40 69.80 47.20 48.10 48.50
3D-LLM Hong et al. (2023b) 35.00 66.00 47.00 69.00 48.00 46.00 48.10
LEO Huang et al. (2023) 46.80 64.10 47.00 60.80 44.20 54.30 52.90
LM4Vision Pang et al. (2023) 34.27 67.05 48.17 68.34 43.87 45.64 48.10
SceneVerseJia et al. (2025) - - - - - - 49.90
SIG3D Man et al. (2024) 35.60 67.20 48.50 71.40 49.10 45.80 52.60
Spartun3D-LLM Zhang et al. (2024b) 49.40 67.30 47.10 63.40 45.40 56.60 54.90
Human 88.53 93.84 88.44 95.27 87.22 88.57 90.06
Table 5: Performance comparison of existing models on SQA3D datasets. The question types include "What," "Is," "How," "Can," "Which," and "Others," with the "Avg" column representing the average performance across all types. The metric used is accuracy, and higher values are more desirable.
Figure 4: Illustration of the broader 3D SQA pipeline: scenes are represented as images or point clouds, questions as a combination of text and visual inputs, and locations as either textual descriptions or coordinates. Features are extracted using task-specific or pretrained encoders, fused in a dedicated module, and passed through an answer prediction head, typically an MLP. Recent methods integrate LVLMs into encoders and fusion modules, enabling zero-shot learning capabilities.
Task Example Question
Object Identification What is the object next to the red chair in the room?
Spatial Reasoning Where is the table located relative to the sofa?
Attribute Querying What is the color of the sphere on the shelf?
Object Counting How many chairs are there in the room?
Attribute Comparison Which is taller, the lamp or the bookshelf?
Multi-hop Reasoning Find the green bottle in the kitchen. What is on the shelf above it?
Navigation Guide the agent to the bedroom and locate the bedside table.
Robotic Manipulation Pick up the blue block and place it on the red cube.
Object Affordance What can be done with the knife on the counter?
Functional Reasoning How would you use the tools in the box to fix the broken chair?
Multi-round Dialogue User: Where is the TV?
Model: It is in the living room on the wall.
User: What is under the TV?
Planning Plan a sequence of actions to make a cup of tea using objects in the kitchen.
Task Decomposition Break down the task of assembling a desk into individual steps.
Table 6: Examples of 3D SQA tasks, identified by their objectives, along with representative example questions. The tasks cover a range of capabilities, including object identification, spatial reasoning, attribute querying, multi-hop reasoning, and planning. These tasks demonstrate the diverse applications and challenges addressed in 3D SQA, requiring models to integrate spatial, semantic, and task-specific understanding.